### v1

@@ Abstract
Image denoising is essential for removing noise in images caused by electric device malfunctions or other factors during image acquisition. It helps preserve image quality and interpretation. Many convolutional autoencoder algorithms have proven effective in image denoising. Owing to their promising efficiency, quantum computers have gained popularity. This study introduces a quantum convolutional autoencoder (QCAE) method for improved image denoising. This method was developed by substituting the representative latent space of the autoencoder with a quantum circuit. To enhance efficiency, we leveraged the advantages of the quantum approximate optimization algorithm (QAOA)-incorporated parameter-shift rule to identify an optimized cost function, facilitating effective learning from data and gradient computation on an actual quantum computer. The proposed QCAE method outperformed its classical counterpart as it exhibited lower training loss and a higher structural similarity index (SSIM) value. QCAE also outperformed its classical counterpart in denoising the MNIST dataset by up to 40% in terms of SSIM value, confirming its enhanced capabilities in real-world applications. Evaluation of QAOA performance across different circuit configurations and layer variations showed that our technique outperformed other circuit designs by 25% on average.
keywords:
[…]
@@ 1 Introduction
Images acquired from image sensors frequently exhibit noise caused by hardware malfunctions and various environmental factors. In this context, noise denotes unpredictable fluctuations in brightness, intensity, or color information within a digital image [11, 33]. Image denoising is pivotal for eliminating these distortions from a given input image, aiming to generate a more refined and clearer image [12, 21]. The denoising process is essential for accurate image analysis and benefits both human interpretation and machine learning (ML)-based assessments [46]. ML-based methods comprise supervised, semisupervized, and unsupervised learning methods [26].
@@
An autoencoder (AE) stands out as an unsupervised learning model designed for input reconstruction using a neural network (NN) model. This class of models demonstrates notable versatility, particularly in the group of learning representations [2]. AEs encompass numerous successful variants, including the variational autoencoder (VAE), convolutional autoencoder (CAE), and denoising autoencoder, as discussed in [37]. They operate by mapping the input to a lower-dimensional space (latent space). These encoding and decoding steps are subsequently used to accurately reconstruct the original image. However, the latent space in AE encounters challenges in accurately capturing complex, nonlinear relationships among data points. Moreover, it prioritizes common data points, leaving gaps in the latent space for less frequent data samples and limiting diversity and usability for tasks such as regeneration.
@@
Recently, the popularity of quantum computers has surged among researchers, driven by their unique computational capabilities, such as superposition. Superposition allows quantum circuits to represent multiple states simultaneously, enabling them to model the complex, nonlinear relationships that classical systems struggle with. This capability can lead to more accurate and efficient data compression than classical AE. However, quantum computers remain suboptimal for production applications [6]. The existing quantum computing (QC) landscape is in the noisy intermediate-scale quantum (NISQ) era [5], indicating that current quantum systems exhibit limitations in terms of the number of qubits, circuit depth, and high noise levels. Because of these factors, variational quantum algorithms (VQAs) have emerged as an ideal choice for implementation, primarily due to their lower qubit requirements and reduced circuit depth. VQAs leverage classical computation to optimize the algorithm parameters effectively [18]. In addition, they exhibit robustness against image noise - a major problem for other types of quantum algorithms. The quantum approximate optimization algorithm (QAOA), a leading VQA, tackles combinatorial optimization by transforming the problem into finding the ground state of its corresponding Hamiltonian.
@@
Due to exponentially increasing computational power and the ability to learn data representations in higher-dimensional spaces, our primary objective in this study is to integrate CAE and QAOA to create an advanced latent space for execution within QC. We enhanced our model by incorporating a trainable quantum variational circuit to develop a hybrid architecture termed a quantum convolutional autoencoder (QCAE). The QCAE was specifically designed for the task of MNIST image denoising. To showcase its efficacy, we conducted a comparative analysis of its performance against its classical counterpart, the classical convolutional autoencoder (CCAE). The contributions of this study are as follows:
  • We present an approach that combines a convolutional autoencoder network with quantum circuits to tackle image denoising efficiently. This integration provides the benefit of using a reduced number of qubits, rendering it particularly well-suited for current NISQ devices.
  • We leverage the QAOA to replace the classical latent space, elevating the input into a higher-dimensional space. This transformation enhances the representation of data, leading to improved accuracy and effectiveness, as QAOA is used to find the minimum objective function (cost function).
  • We leverage an advanced parameter-shift rule (PSR) technique to optimize the QAOA parameters finely, resulting in a substantial improvement in the effectiveness of image denoising through this dual parameter shifting strategy.
@@
  • We rigorously conducted denoising experiments using the MNIST dataset, exploring diverse image noise factors across both quantum simulation and machine scenarios. This comprehensive study aims to refine and advance denoising techniques applied to a widely used MNIST dataset.
@@
The rest of the paper is organized as follows. Section 2 describes both classical and quantum algorithms for image denoising and classification. Section 3 reviews the associated background to understand this paper. In Section 4, we delve into the intricate details of our proposed methodology, leveraging QAOA in conjunction with PSR optimization. Section 5 details the specifics of the dataset and environments employed in our study. In Section 6, we present a comprehensive showcase of our superior performance in image denoising. Finally, Section 7 concludes the paper.
@@ 2 Related works
In the field of classical image denoising, deep learning serves as the main building block, harnessing the power of neural networks (NN) to effectively reduce noise and enhance image quality. Gondara et al. [15] demonstrated that denoising AE, constructed with convolutional layers and applied to small sample sizes, proves to be an effective solution for efficiently denoising medical images. The experiment reveals that even the most basic networks can successfully reconstruct images, even in the presence of corruption levels so severe that the distinction between noise and signal becomes imperceptible to the human eye. Venkataraman et al. [45] employed a straightforward AE network, constructing various architectures for the model. They systematically compared the results to determine the most suitable architecture for image denoising. Paul et al. [32] introduced an efficient denoising method for hyperspectral images (HSI), targeting various noise patterns and their combinations. Using a dual-branch deep neural network based on wavelet-transformed bands, the first branch uses convolutional layers with residual learning for local and global noise feature extraction. The second branch contains an AE layer with subpixel sampling for prominent noise feature extraction. Schneider et al. [39] implemented a CAE, a VAE, and an adversarial autoencoder on two distinct publicly available datasets. Subsequently, we conducted a comprehensive comparison of their anomaly-detection performance. In [38], the author extensively reviewed medical denoising techniques for images, emphasizing their significant importance in medical applications, particularly in disease diagnosis.
@@ 3.1 Quantum computing (QC)
QC harnesses the extraordinary principles of quantum physics, including superposition, entanglement, and quantum measurement, to accelerate data processing capabilities exponentially [40]. This section describes the fundamental and pivotal elements that empower quantum computers.
@@ 3.1.1 Quantum bit (Qubit)
The operational architecture of quantum computers differs fundamentally from that of classical computers. Quantum computers use quantum bits (qubits) as operational units, unlike classical computers, which rely on binary digits existing in either states 0 or 1. A qubit can exist simultaneously in a superposition of states 0, 1, and a linear combination of both 0 and 1 [28]. A qubit is defined as a linear combination of $|0\rangle$ and $|1\rangle$ which are eigen states of SU(2) Hermitian Hamiltonian as follows:
$|0\rangle=\begin{bmatrix}1\\ 0\end{bmatrix},\;|1\rangle=\begin{bmatrix}0\\ 1\end{bmatrix},\;|\psi\rangle=\alpha|0\rangle+\beta|1\rangle=\begin{bmatrix}\alpha\\ \beta\end{bmatrix},$ (1)
@@ 3.1.2 Quantum data manipulation
$H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\quad CNOT=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{bmatrix},\end{array}$ (2)
$R_{X}(\theta)=\begin{bmatrix}cos\Bigl{(}\frac{\theta}{2}\Bigl{)}&-i\cdot sin\Bigl{(}\frac{\theta}{2}\Bigl{)}\\ \\ -i\cdot sin\Bigl{(}\frac{\theta}{2}\Bigl{)}&cos\Bigl{(}\frac{\theta}{2}\Bigl{)}\end{bmatrix},$ (3)
$R_{Y}(\theta)=\begin{bmatrix}cos\Bigl{(}\frac{\theta}{2}\Bigl{)}&-sin\Bigl{(}\frac{\theta}{2}\Bigl{)}\\ \\ sin\Bigl{(}\frac{\theta}{2}\Bigl{)}&cos\Bigl{(}\frac{\theta}{2}\Bigl{)}\end{bmatrix},$ (4)
$R_{Z}(\theta)=\begin{bmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{bmatrix}.$ (5)
@@ 3.2 Classical data encoding
Numerous QML algorithms require converting classical data into quantum states within quantum computers, a process known as data encoding [9, 48]. In other words, the dataset is initially translated from the subject data domain $D$ to the Hilbert space $H$ through a designated feature mapping process $f:D\rightarrow H$. One of the encoding techniques, specifically angle encoding or qubit encoding, is widely renowned for its ability to leverage a minimal number of qubits that correspond to the size of the input vector [30]. The angle encoding scheme demonstrates remarkable efficiency as it necessitates the rotation of only a single qubit as depicted in Fig. 1.
@@
In this technique, each data value $x$ undergoes an initial normalization process, mapping it to the range $[0,2\pi]$. Subsequently, it is encoded using single-qubit rotation gates $R_{X}$, $R_{Y}$, and $R_{Z}$. These rotation gates dynamically determine the rotation angle $\theta$ based on the corresponding data value $x$. The predominant technique involves applying the feature map $x\rightarrow cos\Bigl{(}\frac{x}{2}\Bigl{)}|0\rangle+sin\Bigl{(}\frac{x}{2}\Bigl{)}|1\rangle$ using the $R_{Y}$ as follows:
$R_{Y}(x)|0\rangle&=\begin{bmatrix}cos\Bigl{(}\frac{x}{2}\Bigl{)}&-sin\Bigl{(}\frac{x}{2}\Bigl{)}\\ \\ sin\Bigl{(}\frac{x}{2}\Bigl{)}&cos\Bigl{(}\frac{x}{2}\Bigl{)}\end{bmatrix}\begin{bmatrix}1\\ \\ 0\end{bmatrix}\\ \\ &=cos\Bigl{(}\frac{x}{2}\Bigl{)}|0\rangle+sin\Bigl{(}\frac{x}{2}\Bigl{)}|1\rangle.\end{array}$ (6)
@@ 3.3 Parameterized quantum circuit (PQC)
A PQC provides a tangible means to implement algorithms and showcase quantum supremacy in the NISQ era [4, 19]. Conceptually similar to a classical NN, it features trainable parameters within circuits comprising fixed gates, such as CNOT, and adjustable gates like rotation gates. It typically requires a low number of qubits and circuit depth. However, even with a limited number of qubits and circuit depth, certain classes of PQCs exhibit the ability to generate highly non-trivial outputs. Similar to classical NN, a higher circuit depth repeated in a PQC provides increased flexibility to the circuit, leading to superior results in various applications.
Consider a trainable unitary $U_{\theta}$ operating on an $n$-qubit state applied to a reference state $|\phi\rangle$, where the trained variable $|\phi_{\theta}\rangle=U_{\theta}|\phi\rangle$ generally equals to $|0\rangle^{\otimes n}$. Nevertheless, in the context of an $n$-qubit encoding, which represents a $2^{n}\times 2^{n}$ unitary mapping, the number of input qubits equals the number of output qubits. For the 8-qubit circuit, we encounter a total of $5362^{16}=65,$ 536 parameters, which proves to be an excessively high number for efficient optimization. Therefore, to enhance efficiency, we constrained gate shapes to have fewer parameters using small, efficient local gates, such as CNOT and rotation gates, as the fundamental building blocks of the overall unitary operation.
@@ 3.4 Classical convolutional autoencoder (CCAE)
An autoencoder (AE) is an unsupervised NN comprising an encoder $E$ and a decoder $D$, coined in the 1980s [3]. It simply reconstructs input image data within its hidden layers, using the input as the expected output. The CCAE enhances this by substituting fully connected layers with convolutional layers, optimizing it for spatial data tasks, and ensuring superior feature extraction and representation learning. Like a typical AE, the CCAE maintains input and output layer sizes but transforms the decoder with transposed convolutional layers for enhanced decoding capabilities [50]. Fig. 2 illustrates the architecture of the CCAE, showcasing the process of deconstructing and reconstructing input image data. Through this transformation, the encoded data traverses the latent space $Z$, enabling the model to learn and focus specifically on crucial components of the input image data.
Figure 2: Architecture of classical convolutional autoencoder: The core design of a convolutional autoencoder consists of classical encoder and decoder components, each featuring crucial convolutional layers. These layers play a pivotal role in transforming the input image data $x$ into into a reconstructed representation $\hat{X}$ through the latent space $Z$.
Given an input image data $x$, the process of CCAE is simply defined as follows:
@@
where the training loss can be defined as follows:

### v2

@@ Abstract
Image denoising is essential for removing noise in images caused by electric device malfunctions or other factors during image acquisition. It ensures the preservation of image quality and accurate interpretation. Many convolutional autoencoder algorithms have proven effective in image denoising. Owing to their promising efficiency, quantum computers have gained popularity. This paper proposes a method, the quantum convolutional autoencoder (QCAE), which enhances traditional convolutional autoencoders by replacing their latent space with a quantum counterpart implemented via a QAOA-inspired ansatz circuit. To enhance efficiency, we leveraged the advantages of the quantum approximate optimization algorithm (QAOA)-incorporated parameter-shift rule to identify an optimized cost function, facilitating effective learning from data and gradient computation on an actual quantum computer. The proposed QCAE method outperformed its classical counterpart as it exhibited lower training loss and a higher structural similarity index (SSIM) value. QCAE also outperformed its classical counterpart in denoising the MNIST dataset by up to 40% in terms of SSIM value, confirming its enhanced capabilities in real-world applications. Evaluation of QAOA performance across different circuit configurations and layer variations showed that our technique outperformed other circuit designs by 25% on average.
: Mach. Learn.: Sci.
Keywords: Quantum computing, Quantum convolutional autoencoder, Variational quantum algorithm, Quantum machine learning
@@ 1 Introduction
Images acquired from image sensors frequently exhibit noise caused by hardware malfunctions and various environmental factors. In this context, noise denotes unpredictable fluctuations in brightness, intensity, or color information within a digital image [1, 2]. Image denoising is pivotal for eliminating these distortions from a given input image, aiming to generate a more refined and clearer image [3, 4]. The denoising process is essential for accurate image analysis and benefits both human interpretation and machine learning (ML)-based assessments [5].
@@
An autoencoder (AE) is an unsupervised neural network model designed for reconstructing its input through a learned latent representation. This class of models demonstrates notable versatility, particularly in the group of learning representations [6]. AEs encompass numerous successful variants, including the variational autoencoder (VAE), convolutional autoencoder (CAE), and denoising autoencoder, as discussed in [7]. They operate by mapping the input to a lower-dimensional space (latent space). These encoding and decoding steps are subsequently used to accurately reconstruct the original image. However, the latent space in AE encounters challenges in accurately capturing complex, nonlinear relationships among data points. Moreover, it often prioritizes frequent data patterns, resulting in sparse or poorly structured latent representations that underrepresent rare samples. This can reduce effectiveness in generative or reconstruction tasks [8].
@@
Recently, the popularity of quantum computers has surged among researchers, driven by their unique computational capabilities, such as superposition. Quantum circuits leverage principles such as superposition, entanglement, and interference, allowing them to represent and manipulate complex, high-dimensional states that are often intractable for classical systems. This capability can lead to more accurate and efficient data compression than classical AE. However, quantum computers remain suboptimal for production applications [9]. The existing quantum computing (QC) landscape is in the noisy intermediate-scale quantum (NISQ) era [10], indicating that current quantum systems exhibit limitations in terms of the number of qubits, circuit depth, and high noise levels. Because of these factors, variational quantum algorithms (VQAs) have emerged as an ideal choice for implementation, primarily due to their lower qubit requirements and reduced circuit depth. VQAs leverage classical computation to optimize the algorithm parameters effectively [11]. In addition, they exhibit robustness against image noise - a major problem for other types of quantum algorithms. We adopt a QAOA-inspired ansatz circuit, leveraging its parameterized structure to construct a quantum latent space representation, rather than applying QAOA for solving an optimization task. These quantum principles, superposition, entanglement, and interference, allow the circuit to represent and process complex, high-dimensional patterns in data. In our approach, this variational structure enables effective learning and denoising of noisy images through gradient-based optimization, rather than through solving a combinatorial objective.
@@
Due to exponentially increasing computational power and the ability to learn data representations in higher-dimensional spaces, our primary objective in this study is to integrate CAE and QAOA to create an advanced latent space for execution within QC. We enhanced our model by incorporating a trainable quantum variational circuit to develop a hybrid architecture termed a quantum convolutional autoencoder (QCAE). The QCAE was specifically designed for the task of MNIST [12] image denoising. To showcase its efficacy, we conducted a comparative analysis of its performance against its classical counterpart, the classical convolutional autoencoder (CCAE). The contributions of this study are as follows:
  • We present an approach where the convolutional autoencoder ensures efficient feature extraction, while the QAOA-inspired ansatz circuit contributes the advantage of reduced qubit requirements and improved suitability for NISQ devices.
  • We leverage a QAOA-inspired ansatz circuit to replace the classical latent space, enabling the representation of input data in a higher-dimensional quantum space for improved feature extraction and denoising.
  • We apply the parameter-shift rule (PSR) to efficiently and precisely compute gradients of the QAOA-inspired ansatz circuit, enabling stable optimization within hybrid quantum-classical training.
@@
  • We rigorously conducted denoising experiments using the MNIST dataset, exploring diverse image noise (Gaussian noise) factors across both quantum simulation and machine scenarios.
@@
The rest of the paper is organized as follows. Section 2 describes both classical and quantum algorithms for image denoising and classification. Section 3 reviews the associated background to understand this paper. In Section 4, we delve into the details of our proposed methodology, leveraging QAOA in conjunction with PSR optimization. Section 5 details the specifics of the dataset and environments employed in our study. In Section 6, we present a comprehensive showcase of our superior performance in image denoising. Finally, Section 7 concludes the paper.
@@ 2 Related works
In the field of classical image denoising, deep learning serves as the main building block, harnessing the power of neural networks (NN) to effectively reduce noise and enhance image quality.
In the field of classical image denoising, deep learning serves as the main building block, harnessing the power of neural networks to effectively reduce noise and enhance image quality [3, 4]. Autoencoders and their variants, including convolutional autoencoders, have been widely applied to natural images and benchmark datasets, demonstrating strong denoising performance [13]. These methods form the baseline context against which hybrid quantum–classical approaches, such as our QCAE, can be evaluated.
@@ 3.1 Quantum computing (QC)
Quantum computing leverages fundamental principles such as superposition, entanglement, and quantum interference to explore new computational paradigms, with the potential for advantages in specific tasks over classical systems  [19].
@@ 3.1.1 Quantum bit (Qubit)
The operational architecture of quantum computers differs fundamentally from that of classical computers. Quantum computers use quantum bits (qubits) as operational units, unlike classical computers, which rely on binary digits existing in either state 0 or 1. A qubit can exist simultaneously in a superposition of states 0, 1, and a linear combination of both 0 and 1 [20]. The basis states of $|0\rangle$ and $|1\rangle$ represent the computational basis states of a qubit, forming an orthonormal basis in a two-dimensional Hilbert space. Any pure qubit state can be expressed as a linear combination of these basis states, defined as follows:
$\left|0\right\rangle=\left[\!\begin{array}[]{c}1\\[2.0pt] 0\end{array}\!\right],\quad\left|1\right\rangle=\left[\!\begin{array}[]{c}0\\[2.0pt] 1\end{array}\!\right],\quad\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle=\left[\!\begin{array}[]{c}\alpha\\[2.0pt] \beta\end{array}\!\right].$ (1)
@@ 3.1.2 Quantum data manipulation
$\begin{array}[]{ll}H=\frac{1}{\sqrt{2}}\left[\!\begin{array}[]{cc}1&1\\[2.0pt] 1&-1\end{array}\!\right],\quad\mathrm{CNOT}=\left[\!\begin{array}[]{cccc}1&0&0&0\\[2.0pt] 0&1&0&0\\[2.0pt] 0&0&0&1\\[2.0pt] 0&0&1&0\end{array}\!\right].\end{array}$ (2)
$R_{X}(\theta)=\left[\!\begin{array}[]{cc}\cos\!\Bigl(\frac{\theta}{2}\Bigr)&-i\,\sin\!\Bigl(\frac{\theta}{2}\Bigr)\\[4.0pt] -i\,\sin\!\Bigl(\frac{\theta}{2}\Bigr)&\cos\!\Bigl(\frac{\theta}{2}\Bigr)\end{array}\!\right].$ (3)
$R_{Y}(\theta)=\left[\!\begin{array}[]{cc}\cos\!\Bigl(\frac{\theta}{2}\Bigr)&-\sin\!\Bigl(\frac{\theta}{2}\Bigr)\\[4.0pt] \sin\!\Bigl(\frac{\theta}{2}\Bigr)&\cos\!\Bigl(\frac{\theta}{2}\Bigr)\end{array}\!\right].$ (4)
$R_{Z}(\theta)=\left[\!\begin{array}[]{cc}e^{-i\theta/2}&0\\[4.0pt] 0&e^{i\theta/2}\end{array}\!\right].$ (5)
@@ 3.2 Classical data encoding
Numerous QML algorithms require converting classical data into quantum states within quantum computers, a process known as data encoding [22, 23]. In other words, the dataset is initially translated from the subject data domain $D$ to the Hilbert space $H$ through a designated feature mapping process $f:D\rightarrow H$. One of the encoding techniques, specifically angle encoding or qubit encoding, is widely renowned for its ability to leverage a minimal number of qubits that correspond to the size of the input vector [24]. The angle encoding scheme demonstrates remarkable efficiency as it necessitates the rotation of only a single qubit, as depicted in Fig. 1.
@@
In this technique, each data value $x$ undergoes an initial normalization process, mapping it to the range $0,2\pi]$. Subsequently, it is encoded using single-qubit rotation gates $R_{X}$, $R_{Y}$, and $R_{Z}$. These rotation gates dynamically determine the rotation angle $\theta$ based on the corresponding data value $x$. The predominant technique involves applying the feature map $x\rightarrow cos\Bigl(\frac{x}{2}\Bigl)|0\rangle+sin\Bigl(\frac{x}{2}\Bigl)|1\rangle$ using the $R_{Y}$ as follows:
$\begin{array}[]{ll}R_{Y}(x)\left|0\right\rangle&=\left[\!\begin{array}[]{cc}\cos\!\Bigl(\frac{x}{2}\Bigr)&-\sin\!\Bigl(\frac{x}{2}\Bigr)\\[4.0pt] \sin\!\Bigl(\frac{x}{2}\Bigr)&\cos\!\Bigl(\frac{x}{2}\Bigr)\end{array}\!\right]\left[\!\begin{array}[]{c}1\\[4.0pt] 0\end{array}\!\right]\\[10.0pt] &=\cos\!\Bigl(\frac{x}{2}\Bigr)\left|0\right\rangle+\sin\!\Bigl(\frac{x}{2}\Bigr)\left|1\right\rangle.\end{array}$ (6)
@@ 3.3 Parameterized quantum circuit (PQC)
A PQC provides a practical framework for implementing variational algorithms on NISQ devices, enabling hybrid quantum-classical computations with limited qubits and circuit depth [25, 26]. Conceptually similar to a classical NN, it features trainable parameters within circuits comprising fixed gates, such as CNOT, and adjustable gates like rotation gates. It typically requires a low number of qubits and circuit depth. However, even with a limited number of qubits and circuit depth, certain classes of PQCs exhibit the ability to generate highly non-trivial outputs. Similar to classical NN, a higher circuit depth repeated in a PQC provides increased flexibility to the circuit, leading to superior results in various applications.
Consider a trainable unitary $U_{\theta}$ operating on an $n$-qubit state applied to a reference state $|\phi\rangle$, where the trained variable $|\phi_{\theta}\rangle=U_{\theta}|\phi\rangle$ generally equals to $|0\rangle^{\otimes n}$. Nevertheless, in the context of an $n$-qubit encoding, which represents a $2^{n}\times 2^{n}$ unitary mapping, the number of input qubits equals the number of output qubits. For the 8-qubit circuit, we encounter a total of $2^{16}=65,536$ parameters, which proves to be an excessively high number for efficient optimization. Therefore, to enhance efficiency, we constrained gate shapes to have fewer parameters using small, efficient gates such as single-qubit rotations and entangling gates like CNOT as the fundamental building blocks of the overall unitary operation.
@@ 3.4 Classical convolutional autoencoder (CCAE)
An autoencoder (AE) is an unsupervised NN comprising an encoder $E$ and a decoder $D$, coined in the 1980s [27]. It reconstructs input image data within its hidden layers, using the input as the expected output. The CCAE enhances this by substituting fully connected layers with convolutional layers, optimizing it for spatial data tasks, and ensuring superior feature extraction and representation learning. Like a typical AE, the CCAE maintains input and output layer sizes but transforms the decoder with transposed convolutional layers for enhanced decoding capabilities [28]. Fig. 2 illustrates the architecture of the CCAE, showcasing the process of deconstructing and reconstructing input image data. Through this transformation, the encoded data traverses the latent space $Z$, enabling the model to learn and focus specifically on crucial components of the input image data.
Figure 2: Architecture of classical convolutional autoencoder: The core design of a convolutional autoencoder consists of classical encoder and decoder components, each featuring crucial convolutional layers. These layers play a pivotal role in transforming the input image data $x$ into a reconstructed representation $\hat{X}$ through the latent space $Z$.
Given an input image data $x$, the process of CCAE is defined as follows:
@@
the most commonly used training loss for autoencoders is the L2 reconstruction loss, defined as: